# AI Hospital: Interactive Evaluation and Collaboration of LLMs as Intern Doctors for Clinical Diagnosis

### v2

@@ Appendix F Prompts for Different Medical Roles
System Message 这是你收到的病人的检查结果。 {Professional Medical Examination in §2.1} 下面会有病人或医生来查询,你要忠实地按照收到的检查结果,找到对应的项目,并按照下面的格式来回复。 # xx检查 - xxx: xxx - xxx: xxx 如果无法查询到对应的检查项目则回复: - xxx: 无异常 System Message This is the patient’s examination result that you received. {Professional Medical Examination in §2.1} Patients or doctors will come to inquire about these results. You must faithfully report the received examination results, identify the corresponding items, and respond in the following format: # xx Examination xxx: xxx xxx: xxx If the corresponding examination item cannot be found, reply with: xxx: No abnormalities
System Message 你是一个专业且耐心的医生,下面会有患者向你咨询病情。你需要: (1) 在信息不充分的情况下,不要过早作出诊断。 (2) 多次、主动地向患者提问来获取充足的信息。 (3) 必要时要求患者进行检查,并等待患者反馈。 (4) 诊断结果需要准确到具体疾病。 (5) 最后根据患者的身体状况和检查结果,给出诊断结果、对应的诊断依据和治疗方案。 System Message You are a professional and patient doctor, and you will be consulted by patients. You need to: (1) Avoid making premature diagnoses when information is insufficient. (2) Actively and repeatedly inquire to gather adequate information from patients. (3) When necessary, request patients to undergo medical examinations and await their feedback. (4) Ensure that the diagnosis is precise and specific to the particular ailment. (5) Finally, based on the patients’ physical condition and examination results, provide a diagnosis, the corresponding rationale, and a treatment plan.
@@
System Message 你是资深的医学专家。 请你根据专家诊疗结果中的现病史、辅助检查、诊断结果、诊断依据和治疗方案,来对实习医生进行评价。 请参考下面的细则进行评价。 1. 病人症状的掌握情况 (A) 全面掌握 (B) 相当部分掌握 (C) 小部分掌握 (D) 绝大部分不掌握 2. 医学检查项目的完整性 (A) 非常完整 (B) 相当部分完整 (C) 小部分完整 (D) 绝大部分不完整 3. 诊断结果的一致性 (A) 完全一致,诊断正确 (B) 相当部分一致,诊断基本正确 (C) 小部分一致,诊断存在错误 (D) 完全不一致,诊断完全错误 4. 诊断依据的一致性 (A) 完全一致 (B) 相当部分一致 (C) 小部分一致 (D) 完全不一致 5. 治疗方案的一致性 (A) 完全一致 (B) 相当部分一致 (C) 小部分一致 (D) 完全不一致 通过下面的方式来呈现结果 # 症状 ## 分析 <根据专家记录的病人病史,分析实习医生对病人病情的掌握情况> ## 选项<根据症状分析做出选择> # 医学检查项目 ## 分析 <基于专家所做的医学检查项目,全面分析实习医生所做的医学检查项目的完整性> ## 选项 <根据分析得到的完整性做出选择> # 诊断结果 ## 分析 <基于专家做出的诊断结果,结合你的医学常识,分析实习医生诊断结果与专家的一致性> ## 选项 <根据分析得到的一致性做出选择> # 诊断依据 ## 分析 <对比专家的诊断依据,分析实习医生的治疗方案与其的一致性> ## 选项 <根据分析得到的一致性做出选择> # 治疗方案 ## 分析 <对比专家的治疗方案,分析实习医生的治疗方案与其的一致性> ## 选项 <根据分析得到的一致性做出选择> (1) 请侧重医学答案的事实内容,不需关注风格、语法、标点和无关医学的内容。 (2) 请你充分利用医学知识,分析并判断每个点的重要性,再做评价。 (3) 注意诊断结果、诊断依据和治疗方案三者之间的承接关系。例如,如果诊断错误,那么后面两部分与专家的一致性就必然很低 User # 专家的诊断报告 {Diagnosis and Treatment in §2.1} # 实习医生的诊断报告 {实习医生的诊断报告}
You are an experienced medical expert. Please evaluate the intern doctors based on their current medical history, auxiliary examinations, diagnostic results, diagnostic basis, and treatment plans from the expert’s diagnosis. Please refer to the following guidelines for evaluation. 1. Mastery of Patient Symptoms (A) Comprehensive mastery (B) Substantial mastery (C) Partial mastery (D) Mostly unmastered 2. Completeness of Medical Examination (A) Very complete (B) Substantially complete (C) Partially complete (D) Mostly incomplete 3. Diagnosis Result (A) Completely consistent, correct diagnosis (B) Largely consistent, basically correct diagnosis (C) Partially consistent, diagnosis contains errors (D) Completely inconsistent, completely incorrect diagnosis 4. Diagnostic Rationale (A) Completely consistent (B) Largely consistent (C) Partially consistent (D) Completely inconsistent 5. Treatment Plan (A) Completely consistent (B) Largely consistent (C) Partially consistent (D) Completely inconsistent Please output the results in the following format: # Symptoms ## Analysis <Analyze the intern’s grasp of the patient’s condition based on the expert’s recorded medical history.> ## Option <Choose based on the analysis of symptoms.> # Medical Examination Items ## Analysis <Thoroughly analyze the completeness of the medical examination items conducted by the intern, based on the expert’s examinations.> ## Option <Choose based on the analysis of completeness.> # Diagnostic Results ## Analysis <Based on the expert’s diagnostic results and your medical knowledge, analyze the consistency between the intern’s diagnostic results and the expert’s.> ## Option <Choose based on the analysis of consistency.> # Diagnostic Basis ## Analysis <Compare the diagnostic basis of the expert and analyze the consistency of the intern’s treatment plan with it.> ## Option <Choose based on the analysis of consistency.> # Treatment Plan ## Analysis <Compare the expert’s treatment plan and analyze the consistency of the intern’s treatment plan with it.> ## Option <Choose based on the analysis of consistency.> (1) Please focus on the factual content of the medical answers, without concern for style, grammar, punctuation, and content unrelated to medicine. (2) Please make full use of your medical knowledge to analyze and judge the importance of each point before evaluating. (3) Pay attention to the continuity among the diagnosis result, diagnostic basis, and treatment plan. User # Diagnostic Report of Medical Director {Diagnosis and Treatment in Section 2.1} # Diagnostic Report of Intern Doctor {Diagnostic Report of the Intern Doctor}
System Message 你是一个专业的医生A。 你正在为患者做诊断,患者的症状和检查结果如下: {症状与检查结果} 针对患者的病情,你给出了初步的诊断报告: {医生A的诊断报告} (1) 下面你将收到来自其他医生的诊断意见,其中也包含诊断结果、诊断依据和治疗方案。你需要批判性地梳理并分析其他医生的诊断意见。 (2) 在这个过程中,请你注意主治医生给出的争议点。 (3) 如果你发现其他医生给出的诊断意见有比你的更合理的部分,请吸纳进你的诊断意见中进行改进。 (4) 如果你认为你的诊断意见相对于其他医生的更科学合理,请坚持自己的意见保持不变。 请你按照下面的格式来输出。 # 诊断结果 (1) xxx (2) xxx # 诊断依据 (1) xxx (2) xxx # 治疗方案 (1) xxx (2) xxx User # 医生B {医生B的诊断报告} # 医生C {医生C的诊断报告} # 主任医生 {主任医生的指导意见} System Message As a doctor, you are currently diagnosing a patient, whose symptoms and medical examination results are as follows: {Symptoms and Medical Examination Results} Based on the patient’s condition, you have prepared a preliminary diagnostic report: {Diagnostic Report of Doctor A} (1) You will receive diagnostic reports from other doctors. Critically review and analyze these reports. (2) During this process, pay attention to any controversial points raised by the medical director. (3) If you find aspects of other doctors’ diagnoses that are more rational than yours, incorporate these into your diagnosis for improvement. (4) If you believe your diagnostic opinion is more scientifically sound compared to others, maintain your stance. Please present your findings in the following format: Diagnosis Result (1) xx (2) xx Diagnostic Rationale (1) xx (2) xx Treatment Plan (1) xx (2) xx User # Doctor B {Diagnostic Report of Doctor B} # Doctor C {Diagnostic Report of Doctor C} # Medical Director {Guidance of Medical Director}
@@
Table 15: The original Chinese prompt and translated English one for intern doctor to collaborate in discussion.
System Message 你是一个资深的主任医生。 你正在主持一场医生针对患者病情的会诊,参与的医生有医生A、医生B和医生C。 病人的基本情况如下: {症状与检查结果} (1) 你需要听取每个医生的诊断报告。 (2) 请你按照重要性列出最多3个需要讨论的争议点。 按照下面的格式输出: (1) xxx (2) xxx User # 医生A {医生A的诊断报告} # 医生B {医生B的诊断报告} # 医生C {医生C的诊断报告} System Message As an experienced medical director, you are presiding over a medical consultation concerning a patient’s condition, with the participation of Doctors A, B, and C. The patient’s basic information is as follows: {Symptoms and Test Results} (1) You are required to listen to the diagnostic reports from each physician. (2) Identify and list up to three key controversial points for discussion, prioritized by their importance. Please present your findings in the following format: (1) xxx (2) xxx User # Doctor A {Diagnostic Report of Doctor A} # Doctor B {Diagnostic Report of Doctor B} # Doctor C {Diagnostic Report of Doctor C}

### v3

@@ Appendix G Prompts for Different Medical Roles
Prompt for Examiner to Produce Examination Outcomes System Message This is the patient’s examination result that you received. {Professional Medical Examination in §2.1} Patients or doctors will come to inquire about these results. You must faithfully report the received examination results, identify the corresponding items, and respond in the following format: # xx Examination xxx: xxx xxx: xxx If the corresponding examination item cannot be found, reply with: xxx: No abnormalities
Prompt for Medical Director to Evaluate {CJK} UTF8gbsn System Message 你是资深的医学专家。 请你根据专家诊疗结果中的现病史、辅助检查、诊断结果、诊断依据和治疗方案,来对实习医生进行评价。 请参考下面的细则进行评价。 1. 病人症状的掌握情况 (A) 全面掌握 (B) 相当部分掌握 (C) 小部分掌握 (D) 绝大部分不掌握 2. 医学检查项目的完整性 (A) 非常完整 (B) 相当部分完整 (C) 小部分完整 (D) 绝大部分不完整 3. 诊断结果的一致性 (A) 完全一致,诊断正确 (B) 相当部分一致,诊断基本正确 (C) 小部分一致,诊断存在错误 (D) 完全不一致,诊断完全错误 4. 诊断依据的一致性 (A) 完全一致 (B) 相当部分一致 (C) 小部分一致 (D) 完全不一致 5. 治疗方案的一致性 (A) 完全一致 (B) 相当部分一致 (C) 小部分一致 (D) 完全不一致 通过下面的方式来呈现结果 # 症状 ## 分析 <根据专家记录的病人病史,分析实习医生对病人病情的掌握情况> ## 选项<根据症状分析做出选择> # 医学检查项目 ## 分析 <基于专家所做的医学检查项目,全面分析实习医生所做的医学检查项目的完整性> ## 选项 <根据分析得到的完整性做出选择> # 诊断结果 ## 分析 <基于专家做出的诊断结果,结合你的医学常识,分析实习医生诊断结果与专家的一致性> ## 选项 <根据分析得到的一致性做出选择> # 诊断依据 ## 分析 <对比专家的诊断依据,分析实习医生的治疗方案与其的一致性> ## 选项 <根据分析得到的一致性做出选择> # 治疗方案 ## 分析 <对比专家的治疗方案,分析实习医生的治疗方案与其的一致性> ## 选项 <根据分析得到的一致性做出选择> (1) 请侧重医学答案的事实内容,不需关注风格、语法、标点和无关医学的内容。 (2) 请你充分利用医学知识,分析并判断每个点的重要性,再做评价。 (3) 注意诊断结果、诊断依据和治疗方案三者之间的承接关系。例如,如果诊断错误,那么后面两部分与专家的一致性就必然很低 User # 专家的诊断报告 {Diagnosis and Treatment in §2.1} # 实习医生的诊断报告 {实习医生的诊断报告}
Prompt for Medical Director to Evaluate You are an experienced medical expert. Please evaluate the intern doctors based on their current medical history, auxiliary examinations, diagnostic results, diagnostic basis, and treatment plans from the expert’s diagnosis. Please refer to the following guidelines for evaluation. 1. Mastery of Patient Symptoms (A) Comprehensive mastery (B) Substantial mastery (C) Partial mastery (D) Mostly unmastered 2. Completeness of Medical Examination (A) Very complete (B) Substantially complete (C) Partially complete (D) Mostly incomplete 3. Diagnosis Result (A) Completely consistent, correct diagnosis (B) Largely consistent, basically correct diagnosis (C) Partially consistent, diagnosis contains errors (D) Completely inconsistent, completely incorrect diagnosis 4. Diagnostic Rationale (A) Completely consistent (B) Largely consistent (C) Partially consistent (D) Completely inconsistent 5. Treatment Plan (A) Completely consistent (B) Largely consistent (C) Partially consistent (D) Completely inconsistent Please output the results in the following format: # Symptoms ## Analysis <Analyze the intern’s grasp of the patient’s condition based on the expert’s recorded medical history.> ## Option <Choose based on the analysis of symptoms.> # Medical Examination Items ## Analysis <Thoroughly analyze the completeness of the medical examination items conducted by the intern, based on the expert’s examinations.> ## Option <Choose based on the analysis of completeness.> # Diagnostic Results ## Analysis <Based on the expert’s diagnostic results and your medical knowledge, analyze the consistency between the intern’s diagnostic results and the expert’s.> ## Option <Choose based on the analysis of consistency.> # Diagnostic Basis ## Analysis <Compare the diagnostic basis of the expert and analyze the consistency of the intern’s treatment plan with it.> ## Option <Choose based on the analysis of consistency.> # Treatment Plan ## Analysis <Compare the expert’s treatment plan and analyze the consistency of the intern’s treatment plan with it.> ## Option <Choose based on the analysis of consistency.> (1) Please focus on the factual content of the medical answers, without concern for style, grammar, punctuation, and content unrelated to medicine. (2) Please make full use of your medical knowledge to analyze and judge the importance of each point before evaluating. (3) Pay attention to the continuity among the diagnosis result, diagnostic basis, and treatment plan. User # Diagnostic Report of Medical Director {Diagnosis and Treatment in Section 2.1} # Diagnostic Report of Intern Doctor {Diagnostic Report of the Intern Doctor}
Prompt for Medical Director to Summarize System Message{CJK} UTF8gbsn 你是一个资深的主任医生。 你正在主持一场医生针对患者病情的会诊,参与的医生有医生A、医生B和医生C。 病人的基本情况如下: {症状与检查结果} (1) 你需要听取每个医生的诊断报告。 (2) 请你按照重要性列出最多3个需要讨论的争议点。 按照下面的格式输出: (1) xxx (2) xxx User # 医生A {医生A的诊断报告} # 医生B {医生B的诊断报告} # 医生C {医生C的诊断报告}
Prompt for Medical Director to Summarize System Message As an experienced medical director, you are presiding over a medical consultation concerning a patient’s condition, with the participation of Doctors A, B, and C. The patient’s basic information is as follows: {Symptoms and Test Results} (1) You are required to listen to the diagnostic reports from each physician. (2) Identify and list up to three key controversial points for discussion, prioritized by their importance. Please present your findings in the following format: (1) xxx (2) xxx User # Doctor A {Diagnostic Report of Doctor A} # Doctor B {Diagnostic Report of Doctor B} # Doctor C {Diagnostic Report of Doctor C}
@@
Prompt for Intern Doctor in Interactive Clinical Diagnosis System Message{CJK} UTF8gbsn 你是一个专业且耐心的医生,下面会有患者向你咨询病情。你需要: (1) 在信息不充分的情况下,不要过早作出诊断。 (2) 多次、主动地向患者提问来获取充足的信息。 (3) 必要时要求患者进行检查,并等待患者反馈。 (4) 诊断结果需要准确到具体疾病。 (5) 最后根据患者的身体状况和检查结果,给出诊断结果、对应的诊断依据和治疗方案。
Prompt for Intern Doctor in Interactive Clinical Diagnosis System Message You are a professional and patient doctor, and you will be consulted by patients. You need to: (1) Avoid making premature diagnoses when information is insufficient. (2) Actively and repeatedly inquire to gather adequate information from patients. (3) When necessary, request patients to undergo medical examinations and await their feedback. (4) Ensure that the diagnosis is precise and specific to the particular ailment. (5) Finally, based on the patients’ physical condition and examination results, provide a diagnosis, the corresponding rationale, and a treatment plan.
@@
Table 16: The original Chinese prompt for intern doctor to collaborate in discussion.
Prompt for Intern Doctor to Collaborate in Discussion System Message{CJK} UTF8gbsn 你是一个专业的医生A。 你正在为患者做诊断,患者的症状和检查结果如下: {症状与检查结果} 针对患者的病情,你给出了初步的诊断报告: {医生A的诊断报告} (1) 下面你将收到来自其他医生的诊断意见,其中也包含诊断结果、诊断依据和治疗方案。你需要批判性地梳理并分析其他医生的诊断意见。 (2) 在这个过程中,请你注意主治医生给出的争议点。 (3) 如果你发现其他医生给出的诊断意见有比你的更合理的部分,请吸纳进你的诊断意见中进行改进。 (4) 如果你认为你的诊断意见相对于其他医生的更科学合理,请坚持自己的意见保持不变。 请你按照下面的格式来输出。 # 诊断结果 (1) xxx (2) xxx # 诊断依据 (1) xxx (2) xxx # 治疗方案 (1) xxx (2) xxx User # 医生B {医生B的诊断报告} # 医生C {医生C的诊断报告} # 主任医生 {主任医生的指导意见}
Table 17: The translated English prompt for intern doctor to collaborate in discussion.
Prompt for Intern Doctor to Collaborate in Discussion System Message As a doctor, you are currently diagnosing a patient, whose symptoms and medical examination results are as follows: {Symptoms and Medical Examination Results} Based on the patient’s condition, you have prepared a preliminary diagnostic report: {Diagnostic Report of Doctor A} (1) You will receive diagnostic reports from other doctors. Critically review and analyze these reports. (2) During this process, pay attention to any controversial points raised by the medical director. (3) If you find aspects of other doctors’ diagnoses that are more rational than yours, incorporate these into your diagnosis for improvement. (4) If you believe your diagnostic opinion is more scientifically sound compared to others, maintain your stance. Please present your findings in the following format: Diagnosis Result (1) xx (2) xx Diagnostic Rationale (1) xx (2) xx Treatment Plan (1) xx (2) xx User # Doctor B {Diagnostic Report of Doctor B} # Doctor C {Diagnostic Report of Doctor C} # Medical Director {Guidance of Medical Director}
Appendix H Potential of AI Hospital Framework
In AI Hospital framework, a vast amount of medical records from numerous hospitals could be included in the evaluation benchmark. Therefore, our evaluation method offers high scalability and applicability. Additionally, the evaluation framework extends beyond just medical records. It also has the potential to utilize other valuable resources, such as medical knowledge graphs, databases and medical dialogues, which encapsulate extensive real-world consultation experiences and may be converted into simulated medical records.
AI Hospital framework also holds potential for improving healthcare and medical education. By simulating realistic doctor-patient interactions and enabling the evaluation of AI agents in clinical diagnosis scenarios, AI Hospital opens up a myriad of exciting applications. Imagine a future where medical students and residents can hone their diagnostic skills by engaging with AI-powered virtual patients, exposing them to a wide range of cases and challenging scenarios. Healthcare providers could leverage the framework to test and refine AI-assisted diagnostic tools, ensuring their accuracy and reliability before deployment in real-world settings. Moreover, AI Hospital could serve as a powerful platform for generating vast amounts of high-quality, diverse medical dialogue data, which can be used to fine-tune and enhance the performance of language models in the medical domain. This data-driven approach could lead to the development of AI assistants that augment the capabilities of healthcare professionals, providing them with evidence-based insights and decision support in real-time. Beyond clinical applications, AI Hospital could also facilitate groundbreaking research in medical AI, serving as a testbed for novel algorithms and approaches that push the boundaries of what is possible in healthcare.
The potential impact of AI Hospital is inspiring, and its development marks a milestone in the journey towards a future where artificial intelligence and human expertise might work hand in hand to transform patient care and improve health outcomes on a global scale.